OpenStudy (anonymous):

(x-2)180=1080 I forgot how to find x

OpenStudy (anonymous):

X equals 8.

OpenStudy (anonymous):

(x-2)180=1080 180x-360=1080 180x=1080-360 180x=720 x=4